# Attention-Based Multi-Scale Temporal Fusion Network for Uncertain-Mode Fault Diagnosis in Multimode Processes

### v1

@@ Abstract
Fault diagnosis in multimode processes plays a critical role in ensuring the safe operation of industrial systems across multiple modes. It faces a great challenge yet to be addressed – that is, the significant distributional differences among monitoring data from multiple modes make it difficult for the models to extract shared feature representations related to system health conditions. In response to this problem, this paper introduces a novel method called attention-based multi-scale temporal fusion network. The multi-scale depthwise convolution and gated recurrent unit are employed to extract multi-scale contextual local features and long-short-term features. A temporal attention mechanism is designed to focus on critical time points with higher cross-mode shared information, thereby enhancing the accuracy of fault diagnosis. The proposed model is applied to Tennessee Eastman process dataset and three-phase flow facility dataset. The experiments demonstrate that the proposed model achieves superior diagnostic performance and maintains a small model size.
@@ 1 Introduction
With the rapid development of automation, sensor and computer technologies, the scale and complexity of industrial systems have grown significantly. The complex industrial systems pose significant challenges in terms of safety and reliability engineering[1, 2]. The increasing number of components in these systems leads to more frequent and complex interactions, greatly raising the probability of faults [3]. The complexity of the processes and the propagation of faults make fault classification increasingly difficult. The occurrence of faults can result in equipment downtime, production interruptions, and resource waste, leading to significant economic losses and even posing a threat to personnel safety [4]. Therefore, developing effective fault diagnosis methods is essential to ensure the safe and reliable operation of industrial systems.
@@
With the deployment of a large number of sensors, industrial systems collect and store massive amounts of process data, and the exploration of data-driven fault diagnosis methods has become increasingly popular [5, 6]. Data-driven fault diagnosis methods include statistical methods and machine learning methods. The statistical methods for fault diagnosis [7] include principal component analysis (PCA), partial least squares (PLS), independent component analysis (ICA), Fisher discriminant analysis (FDA) [8, 9, 10, 11]. The machine learning methods for fault diagnosis include models such as support vector machine (SVM),k-nearest neighbors (KNN), and Bayesian network (BN) [12, 13, 14]. These methods have limitations when dealing with high-dimensional data and often require significant domain expertise for feature extraction. Compared to traditional machine learning methods, deep learning methods have advantages such as automatic feature extraction and higher accuracy [15]. Currently, deep learning methods applied to fault diagnosis in industrial systems include convolutional neural networks (CNN), stacked autoencoders (SAE), long-short-term memory networks (LSTM), temporal convolutional networks (TCN), and visual transformers (ViT). Wu et al. [16] constructed a fault diagnosis method based on deep convolutional neural networks (DCNN). This model incorporates convolutional layers, pooling layers, dropout and fully connected layers to extract feature representations. Liu et al. [17] integrated the sparse autoencoder with the denoising autoencoder to capture robust, sparse but intrinsic nonlinear features. Huang et al. proposed a fault diagnosis model named CNN-LSTM based on a combination of CNN and LSTM. The CNN layers and the LSTM layers are used to perform feature learning and capture time delay information respectively [18]. Zhou et al. proposed industrial process optimization vision transformer (IPO-ViT) fault diagnosis method based on ViT [19]. The self-attention mechanism in this model is used to perform global feature learning on the complex process signals. Zhu et al. combined the multi-scale feature extraction and the deep bidirectional mechanism to better extract the time series features [20]. These models primarily focus on fault diagnosis under a single operating mode. However, as the environmental conditions, loads, and production schedules change, industrial systems often switch between multiple operating modes. The significant distributional differences among monitoring data from multiple modes pose significant challenges for existing fault diagnosis methods.
@@
Recently, some studies have focused on fault diagnosis in multimode processes. These works primarily address the challenge of distribution differences between the training set (source domain) and the testing set (target domain), which are referred to as cross-domain fault diagnosis. The techniques employed in cross-domain fault diagnosis methods are classified into data extension, learning strategy, and representation learning. Data extension promotes the learning of cross-domain generalizable feature representations by increasing the diversity of training data. The methods for increasing data diversity include data augmentation techniques (e.g., MixUp [21] and data transformations [22]) and data generation via deep neural networks [23, 24]. However, the effectiveness of these methods highly depends on the quality of the extended data. Representation learning mainly focuses on reducing the distribution discrepancy between different domains. These methods are categorized into adversarial-based methods and metric-based methods. In adversarial-based methods, adversarial learning is employed to deceive the discriminator, thereby capturing the domain-invariant features [25, 26, 27, 28, 29, 30]. Metric-based methods explicitly align features from different domains to capture generalizable feature representations. This is typically achieved by incorporating distance metrics into the loss function [31, 32, 33, 34, 35, 36]. Learning strategy exploits general learning methods (e.g., meta-learning [37]) to enhance the model’s generalization capability.
@@
Although these fault diagnosis methods for multimode processes have achieved significant progress, there are still some open research gaps. On the one hand, current research focuses on fault diagnosis in multimode processes by capturing domain-invariant features. However, the varying amount of domain-invariant information contained in different time points is often neglected. Overemphasizing time points with limited domain-invariant information can degrade fault diagnosis performance. On the other hand, the operating mode of the samples is typically inaccessible, which makes methods relying on known operating modes inapplicable.
@@
To address the above problems, a model named attention-based multi-scale temporal fusion network (AMTFN) is proposed. Specifically, the multi-scale depthwise convolution (MSDC) and gated recurrent unit (GRU) are employed to capture multi-scale contextual local features and long-short-term feature representations, respectively. Additionally, a temporal attention mechanism (TAM) is constructed to assign weights to deep features at different time steps, thereby enhancing the importance of specific time features that contain more domain-invariant information. The main contributions of this article are summarized as follows.
1) A significant scenario called uncertain-mode fault diagnosis is studied, where the operating mode of the samples is unavailable.
@@
2) A fault diagnosis model named attention-based multi-scale temporal fusion network is proposed for uncertain-mode fault diagnosis.
3) Experiments on two datasets demonstrate that the AMTFN model can achieve effective fault diagnosis with a compact model size.
The rest of this paper is organized as follows. In Section 2, the problem description, related research areas and preliminary theoretical knowledge of depthwise convolution and GRU are introduced. The proposed method is explained in detail in Section 3. In Section 4, the experiments are conducted on the Tennessee Eastman (TE) process dataset and the three-phase flow dataset to validate the effectiveness of the proposed method. Finally, conclusions are drawn in Section 5.
@@ 2.1 Problem formulation
Due to variations in environmental conditions or production schemes, industrial systems switch between multiple stable modes. Assume that a industrial system is designed to operate in $M$ modes, and the monitoring data collected from these modes is denoted as $D=\{(\bm{x}_{k}^{\text{o}},y_{k})\}_{k=1}^{N}\sim P_{XY}$, where $N$ is the number of samples for the $M$ modes. Here, $\bm{x}_{k}^{\text{o}}\in\mathbb{R}^{v}$ denotes the monitoring data at the $k$-th time point, with $v$ being the dimension of the measured variables, and $y_{k}\in\{1,...,L\}$ denotes the system health condition labels, including one normal category and $L-1$ fault categories. Although it is known that the samples are collected from $M$ modes, the specific mode to which each sample belongs is unknown. In the context of uncertain-mode fault diagnosis, it is assumed that fault data for the $M$ modes have been collected, with each mode containing samples for $L$ health conditions. The goal of uncertain-mode fault diagnosis is to build a model based on the monitoring data from these $M$ modes, such that the health condition of the system in these modes can be effectively identified.
@@ 2.2 Related research areas
There are several research fields related to uncertain-mode fault diagnosis (UMFD), including but not limited to: single mode fault diagnosis (SMFD), domain adaptation-based fault diagnosis (DAFD), and domain generalization-based fault diagnosis (DGFD). The differences between these methods and UMFD are summarized in Table 1.
@@
of the $k$-th sample
SMFD $D^{i}$ $D^{i}$ $o_{k}=i$
DAFD $\{D^{i}\}_{i=1}^{S}\cup\{D^{’j}\}_{j=S+1}^{M}$$\{D^{j}\}_{j=S+1}^{M}$ $o_{k}\in\{1,…,M\}\text{ and }o_{k}=o_{0}$
DGFD $\{D^{i}\}_{i=1}^{S}$$\{D^{j}\}_{j=S+1}^{M}$ $o_{k}\in\{1,…,M\}\text{ and }o_{k}=o_{0}$
UMFD $\{D^{i}\}_{i=1}^{M}$$\{D^{i}\}_{i=1}^{M}$ $o_{k}\in\{1,…,M\}$
@@
In SMFD, both the training and testing datasets are collected from the same mode. UMFD differs from SMFD in two main aspects. First, in UMFD, the operating mode of the sample is uncertain. This means that although it is known that the sample belongs to one of the $M$ operating modes, the specific mode is not identifiable. In the case where the operating modes of the sample are known, a fault diagnosis model can be built for each mode, and the corresponding model can be selected based on the sample’s operating mode. Second, UMFD involves monitoring data from multiple modes in both the training and test datasets. Taking the normal condition and fault 13 of the Tennessee Eastman (TE) process as an example, Fig. 1 presents the data distribution of two monitoring variables under six operational modes. Samples from the same health condition category may present complex data distributions due to mode changes, which puts higher requirements on the feature extraction capability of the fault diagnosis model.
Fig. 1.: Data distribution of monitoring variables under six modes in the TE process.
Cross-domain fault diagnosis aims to address the challenge posed by distribution discrepancies between the training set (source domain) and the testing set (target domain). It can be categorized into domain adaptation- based fault diagnosis (DAFD) and domain generalization-based fault diagnosis (DGFD) based on whether the target domain is accessible during the model training phase. Compared to cross-domain fault diagnosis, UMFD does not involve domain shift between the training and testing sets. However, the operating mode of the samples is inaccessible. The extraction of domain-invariant features is key to cross-domain fault diagnosis. UMFD aims to develop models that can effectively extract shared feature representations related to system health conditions from monitoring data across multiple modes. Therefore, the research on UMFD plays an important role in cross domain fault diagnosis.
@@ 2.4 GRU
Compared to the convolution neural network (CNN), the gated recurrent unit (GRU) have an advantage in processing time-series data. The GRU has fewer parameters and achieve faster convergence [39], and its structure is shown in Fig. 3.
@@
The GRU employs two gates, the reset gate and the update gate, to retain important information and discard irrelevant information. The output of GRU is formulated as follows [40],
@@
where $\sigma$ denotes the Sigmoid activation function. $\bm{W}$ and $\bm{U}$ are learnable weight matrices, which are shared across different time steps.
@@ 3 Proposed method
3.1 Overall structure of AMTFN
The overall structure of AMTFN is shown in Fig. 4. The model consists of three main modules: the feature extractor $E$, the feature fusion module $F$, and the classifier $C$. $E$ is used to capture deep feature representations, after which $F$ performs weighted fusion of features across different time steps. Finally, $C$ is employed to determine the health condition of the system.
Fig. 4.: Overall structure of AMTFN.
@@ 3.2 Feature extractor
The feature extractor consists of MSDC, channel-wise concatenation and GRU. MSDC is used to extract multi-scale contextual local features in parallel using convolution kernels of different sizes. Channel-wise concatenation is then employed to splice the multi-scale local features along the channel dimension, followed by GRU to further extract deep long-term and short-term feature representations.
@@
where $\bm{b}_{k}\in\mathbb{R}^{4v\times w}$, DC denotes the depthwise convolution operation, and the subscript indicates the kernel size. In depthwise convolution, instance normalization is applied to weaken the mode features, and the ReLU function is used as the activation function. Concat represents concatenation along the channel dimension.
Then, the multi-scale features $\bm{b}_{k}$ are fed into the GRU, and the hidden state at the $t$-th time step is formulated as follows,
@@
The output of the feature extractor for the $k$-th sample is formulated as, $\bm{h}_{k}=E\left(\bm{x}_{k}\right)=[\bm{h}_{k(1)},\bm{h}_{k(2)},...,\bm{h}_{k(w)}]$.
@@ 3.3 Feature fusion module
Considering that the amount of domain-invariant information varies across different time points, a temporal attention mechanism (TAM) is incorporated into the feature fusion module to enable the model to focus on features at critical time points. The feature fusion module takes the feature map $\bm{h}_{k}$ as input and infers a temporal attention map $\bm{a}_{k}$ through the TAM, as shown in Fig. 4. The temporal attention map is then used as a weight representing the importance of each time step to obtain the feature fusion result $\bm{f}_{k}$. The overall process of feature fusion can be summarized as follows,
@@
The structure of the temporal attention mechanism is shown in Fig. 5. The relationships between different time steps of the features are utilized to generate the temporal attention map. Average pooling and standard deviation are used to aggregate information along the variable dimension of the feature map. These aggregated features are then fed into their respective fully connected layers to generate the average-pooling attention map $\bm{p}_{1}$ and the standard deviation attention map $\bm{p}_{2}$. The calculations for $\bm{p}_{1}$ and $\bm{p}_{2}$ are formulated as,
@@
where $\sigma_{1}$ denotes the ReLU function, AvgPool denote the average-pooling operations, Std denotes the calculation of standard deviation, $F_{c1}$, $F_{c2}$, $F_{c3}$ and $F_{c4}$ denote the fully connected layer.
@@
3.4 Model implementation and application process
The classifier takes the fused feature $\bm{f}_{k}$ as input and generates the final output through a fully connected layer followed by a softmax layer. The final output $\bm{c}_{k}$ is formulated as,

### v2

# Attention-Based Multiscale Temporal Fusion Network for Uncertain-Mode Fault Diagnosis in Multimode Processes

@@ Abstract
Fault diagnosis in multimode processes plays a critical role in ensuring the safe operation of industrial systems across multiple modes. It faces a great challenge yet to be addressed – that is, the significant distributional differences among monitoring data from multiple modes make it difficult for the models to extract shared feature representations related to system health conditions. In response to this problem, this paper introduces a novel method called attention-based multiscale temporal fusion network. The multiscale depthwise convolution and gated recurrent unit are employed to extract multiscale contextual local features and long-short-term features. Instance normalization is applied to suppress mode-specific information. Furthermore, a temporal attention mechanism is designed to focus on critical time points with higher cross-mode shared information, thereby enhancing the accuracy of fault diagnosis. The proposed model is applied to Tennessee Eastman process dataset and three-phase flow facility dataset. The experiments demonstrate that the proposed model achieves superior diagnostic performance and maintains a small model size. The source code will be available on GitHub at https://github.com/GuangqiangLi/AMTFNet.
@@ 1 Introduction
The significant growth in the scale and complexity of modern industrial systems poses significant challenges to the safety and reliability engineering [1, 2, 3]. The increasing number of components in these systems leads to more frequent and complex interactions, which greatly raises the probability of faults and makes fault classification increasingly difficult [4]. The occurrence of faults can result in equipment downtime, production interruptions, and resource waste[5]. Therefore, it is important to monitor the system health condition in real time through the fault diagnosis model.
@@
With the deployment of a large number of sensors, industrial systems collected massive amounts of monitoring data, and there is a growing interest in data-driven fault diagnosis (DDFD) efforts [6, 7]. Data-driven fault diagnosis methods include statistical methods [8, 9, 10, 11, 12] and machine learning methods [13, 14, 15]. These methods have limitations when dealing with high-dimensional data and often require significant domain expertise for feature extraction. Compared to shallow machine learning methods, deep learning methods have advantages such as automatic feature extraction and higher accuracy [16]. Currently, multiple deep learning methods have been applied to fault diagnosis in industrial systems. Wu et al. [17] used the convolutional neural network (CNN) to extract the feature representation. Liu et al. [18] integrated the sparse autoencoder with the denoising autoencoder to capture robust, sparse but intrinsic nonlinear features. Huang et al. added the long short-term memory (LSTM) network to CNN to capture temporal features [19]. Zhou et al. employed the self-attention mechanism to learn global information [20]. Zhu et al. combined multiscale and bidirectional mechanism for feature extraction [21]. These models primarily focus on fault diagnosis under a single operating mode. However, as the environmental conditions, loads, and production schedules change, industrial systems often switch between multiple operating modes. The mode switching leads to complexity in data distribution, which weakens the performance of existing fault diagnosis methods.
@@
Recently, some studies have focused on fault diagnosis in multimode processes. These works primarily address the challenge of distribution differences between the training data (source domain) and the test data (target domain). This scenario is referred to as cross-domain fault diagnosis. The techniques employed in cross-domain fault diagnosis methods mainly include data extension and representation learning. Data extension promotes the learning of cross-domain generalizable feature representations by improving the variety of training samples. The methods for increasing data diversity include data augmentation techniques (e.g., MixUp [22] and data transformations [23]) and data generation via deep neural networks [24, 25]. However, generating high-quality samples is difficult. Representation learning mainly focuses on reducing the variation between source and target domains. These methods include adversarial-based methods and metric-based methods. In adversarial-based methods, adversarial learning is employed to deceive the discriminator, thereby capturing the domain-invariant features [26, 27, 28, 29, 30, 31]. Metric-based methods explicitly align features from different domains to capture generalizable feature representations. This is typically achieved by incorporating distance metrics into the loss function [32, 33, 34, 35, 36, 37, 38].
@@
Although these fault diagnosis methods for multimode processes have achieved significant progress, some issues are still not well addressed. On the one hand, current research focuses on fault diagnosis in multimode processes by capturing domain-invariant features. However, the varying amount of domain-invariant information contained in different time points is often neglected. Under the influence of faults and control loops, the system undergoes a transient-to-steady evolution. Some time steps may show behaviors inconsistent with the final steady-state impact due to control compensation or system inertia. Uniformly using all time steps may introduce irrelevant or misleading features. Focusing on critical moments helps capture fault responses that are less affected by control, thus containing more domain-invariant features. And overemphasizing time points with limited domain-invariant information can decrease the success of fault diagnosis. On the other hand, the operating mode of the samples is typically inaccessible, which makes methods relying on known operating modes inapplicable.
@@
To address the above problems, a model named attention-based multiscale temporal fusion network (AMTFNet) is proposed. Specifically, the multiscale depthwise convolution (MSDC) and gated recurrent unit (GRU) are employed to capture multiscale contextual local features and long-short-term feature representations, respectively. Additionally, a temporal attention mechanism (TAM) is constructed to assign weights to deep features at different time steps, thereby enhancing the importance of specific time features that contain more domain-invariant information. The primary contributions of this study are outlined below.
@@
(1) An attention-based multi-scale temporal fusion network (AMTFNet) is proposed for uncertain-mode fault diagnosis. The model enables accurate diagnosis of multimode processes without requiring prior knowledge of the specific operating mode.
(2) Under the combined influence of faults and control loops, the system exhibits dynamic behaviors that may be inconsistent with the final steady-state response, potentially introducing misleading features. To address this problem, a temporal attention mechanism is designed to focus on critical time steps with more domain-invariant fault information, thereby enhancing diagnostic performance under diverse operating modes.
(3) Compared to recent advanced methods that do not explicitly address the uncertain-mode scenario considered here, AMTFNet achieves effective fault diagnosis in this novel setting while maintaining a compact model size.
The subsequent sections are arranged as follows. In Section 2, the problem description, related research areas and preliminary theoretical knowledge of depthwise convolution and GRU are introduced. The in-depth explanation of the proposed method is presented in Section 3. In Section 4, the experiments are conducted on the two datasets to demonstrate the performance of the proposed method. The final conclusions are summarized in Section 5.
@@ 2.1 Problem formulation
Industrial systems switch between multiple stable modes as environments and production schedules change. Assume that a industrial system is designed to operate in $M$ modes, and the monitoring data collected from these modes is denoted as $D=\{(\bm{x}_{k}^{\text{o}},y_{k})\}_{k=1}^{N}\sim P_{XY}$, where $N$ denotes the number of samples corresponding to the $M$ modes. Here, $\bm{x}_{k}^{\text{o}}\in\mathbb{R}^{v}$ denotes the monitoring data at the $k$-th time point, with $v$ being the dimension of the measured variables, and $y_{k}\in\{1,...,L\}$ denotes the system health condition labels, including one normal state and $L-1$ fault states. Although it is known that the samples are collected from $M$ modes, the specific mode to which each sample belongs is unknown. In uncertain-mode fault diagnosis, it is assumed that fault data for the $M$ modes have been collected, with each mode containing samples for $L$ health conditions. The goal of uncertain-mode fault diagnosis is to construct a model trained on the monitoring data of these $M$ modes, such that the health condition of the system in these modes can be effectively identified.
@@ 2.2 Related research areas
There are several research fields related to uncertain-mode fault diagnosis (UMFD), including but not limited to: single mode fault diagnosis (SMFD), domain adaptation-based fault diagnosis (DAFD), and domain generalization-based fault diagnosis (DGFD). The comparison between these methods and UMFD is presented in Table 1.
@@
of the $k$-th sample $r_{k}$
SMFD $D^{i}$ $D^{i}$ $r_{k}=i$
DAFD $\{D^{i}\}_{i=1}^{S}\cup\{D^{’j}\}_{j=S+1}^{M}$$\{D^{j}\}_{j=S+1}^{M}$ $r_{k}\in\{1,…,M\}\text{ and }r_{k}=r_{0}$
DGFD $\{D^{i}\}_{i=1}^{S}$$\{D^{j}\}_{j=S+1}^{M}$ $r_{k}\in\{1,…,M\}\text{ and }r_{k}=r_{0}$
UMFD $\{D^{i}\}_{i=1}^{M}$$\{D^{i}\}_{i=1}^{M}$ $r_{k}\in\{1,…,M\}$
@@
In SMFD, both the training and test samples are collected from the same mode. UMFD differs from SMFD in two main aspects. First, in UMFD, the operating mode of the sample is uncertain. This means that although it is known that the sample belongs to one of the $M$ operating modes, the specific mode is not identifiable. In the case where the operating modes of the sample are known, a separate model can be built for each mode, and the corresponding model can be selected based on the sample’s operating mode. Second, UMFD involves monitoring data from multiple modes in both the training and test samples. Taking the two health states of the Tennessee Eastman (TE) process for demonstration, Fig. 1 presents the data distribution of two monitoring variables under six operational modes. Samples from the same health condition category may present complex data distributions due to mode changes, which puts higher requirements on the feature representation capability of the fault diagnosis model.
Fig. 1.: Data distribution of two health conditions under six modes in the TE process.
Cross-domain fault diagnosis focuses on the scenarios where there are distribution discrepancies between training and test samples. It includes domain adaptation-based fault diagnosis (DAFD) and domain generalization-based fault diagnosis (DGFD). The difference between the two is that DAFD relies on unlabeled test samples to train the model. Compared to cross-domain fault diagnosis, UMFD does not involve domain shift between the training and test samples. However, the operating mode of the samples is inaccessible. The extraction of universal features is key to cross-domain fault diagnosis. UMFD aims to develop models that can effectively learn shared representations reflecting system condition from monitoring data across multiple modes. Therefore, the research on UMFD plays an important role in cross domain fault diagnosis.
@@ 2.4 GRU
Compared to the CNN, the gated recurrent unit (GRU) have an advantage in processing time-series data. The GRU has fewer parameters and achieve faster convergence [40], and its structure is shown in Fig. 3.
@@
The GRU employs two gates to retain important information and discard non-essential features. The output of GRU is formulated as follows [41],
@@
where $\sigma$ denotes the Sigmoid activation function. $\bm{W}$ and $\bm{U}$ are learnable weight matrices, which are jointly optimized across all time steps.
@@ 3 Proposed method
3.1 Overall structure of AMTFNet
The overall design of AMTFNet is shown in Fig. 4. The model consists of three primary components: feature extractor $E$, feature fusion module $F$, and classifier $C$. $E$ is used to capture deep feature representations, after which $F$ performs weighted fusion of features across different time steps. Finally, $C$ is employed to determine the health condition of the system.
Fig. 4.: Overall structure of AMTFNet.
@@ 3.2 Feature extractor
The feature extractor consists of MSDC, channel-wise concatenation and GRU. MSDC is used to extract multiscale contextual local features in parallel using convolution kernels of different sizes. Channel-wise concatenation is then employed to splice the multiscale local features along the channel dimension, followed by GRU to further extract deep long-term and short-term feature representations.
@@
where $\bm{b}_{k}\in\mathbb{R}^{4v\times w}$, DC denotes the depthwise convolution operation, and the subscript indicates the kernel size. In depthwise convolution, instance normalization is applied to weaken the mode features, and the activation function adopted is ReLU. Concat represents concatenation in the channel-wise direction.
Then, $\bm{b}_{k}$ is fed into GRU, and the hidden vector at the $t$-th time step is formulated as follows,
@@
The extracted features corresponding to the $k$-th sample is formulated as, $\bm{h}_{k}=E\left(\bm{x}_{k}\right)=[\bm{h}_{k(1)},\bm{h}_{k(2)},...,\bm{h}_{k(w)}]$.
@@ 3.3 Feature fusion module
Considering that the amount of domain-invariant information varies across different time points, a temporal attention mechanism (TAM) is incorporated into the feature fusion module to help the model concentrate on features at key time points. The feature fusion module takes the feature map $\bm{h}_{k}$ as input and infers a temporal attention map $\bm{a}_{k}$ through the TAM, as shown in Fig. 4. $\bm{a}_{k}$ is then used as a weight representing the importance of each time step to obtain the feature fusion result $\bm{f}_{k}$. The overall process of feature fusion can be described as follows,
@@
The structure of the TAM is shown in Fig. 5. The relationships between different time steps of the features are utilized to generate the temporal attention map. Average pooling and standard deviation are used to aggregate information along the variable dimension of the feature map. These aggregated features are then transmitted to their respective fully connected layers to generate the average-pooling attention map $\bm{p}_{1}$ and the standard deviation attention map $\bm{p}_{2}$. The calculations for $\bm{p}_{1}$ and $\bm{p}_{2}$ are formulated as,
@@
where $\sigma_{1}$ denotes the ReLU function, AvgPool denote the average-pooling operations, Std denotes the calculation of standard deviation, $F_{c1}-F_{c4}$ denote the fully connected layers.
@@
3.4 Model output and application process
The classifier takes the fused feature $\bm{f}_{k}$ as input and infers health condition class. The model output $\bm{c}_{k}$ is obtained via a fully connected layer, and is formulated as,